An exhibit hall in a convention center measures 20 by 30 yards. A square area of 9 square yards must be reserved in one corner for registration. How many square yards will remain for exhibit space?
Answer by checkley77(12569)

20*30-9=600-9=591 YDS^2 WILL REMAIN FOR EXHIBIT SPACE.